# Interstellar Medium Modulation of Nonlinear Kinetic Alfvén Morphology in Structured Galactic Environments

### v1

@@ II.3.4 Supernova Remnants
The shell is characterized by a strong peak in electron density. This enhancement is highly asymmetric, exhibiting a sharp density jump at the outer shock front followed by a gradual decline toward the interior due to post-shock relaxation (Reynolds, 1998; Orlando et al., 2007). The blast wave efficiently clears ambient ions from the remnant’s interior, creating a deep density cavity. Additionally, the PWN acts as an efficient source for electron-positron pairs, locally elevating the electron density in the core.
Using the formalism from Sec. II.2, we model this three-component morphology as a superposition of the PWN core, the cavity, and the shell:
@@
Positron Fraction: Since the PWN is a source of relativistic pairs, the positron density follows the same profile as electrons in the PWN (Kennel and Coroniti, 1984; Gaensler and Slane, 2006). We model this central positron source as a Gaussian distribution:
@@
Magnetic Field: In our SNR model, the magnetic field is structured in close analogy to the plasma density. The same forward shock that sweeps up and compresses the gas also compresses and amplifies the magnetic field lines via flux freezing (Vink, 2012). Consequently, the magnetic morphology exhibits a turbulent, high-field shell surrounding a comparatively weaker interior cavity that separates the shell from the central PWN core (Kennel and Coroniti, 1984; Gaensler and Slane, 2006). We model the total SNR magnetic field as the superposition of the shell and the PWN core:
@@
Electron Temperature: While the plasma density and magnetic field exhibit a hollow cavity structure, the temperature profile follows a distinctly different morphology. At the outer shock front, the kinetic energy of the sweeping ejecta is thermalized, creating an immediate jump in temperature (Chevalier, 1977; Vink, 2012). Interior to the shell, the physics of the Sedov-Taylor expansion phase dictates that while the density drops, the pressure remains high to support the expanding shell (Chevalier, 1977). Consequently, the temperature must rise towards the center ($T\propto P/n$) to maintain an approximately isobaric state. Furthermore, the center of the remnant is energized by the relativistic PWN (Gaensler and Slane, 2006; Kennel and Coroniti, 1984).
To model this rising interior thermal structure, we treat the temperature as a superposition of shock heating at the shell and a broad, hot core filling the interior:
@@
II.4 Total ISM Profiles and Equilibrium Ion Density
@@
While the NE2001 (Cordes and Lazio, 2002) and JF12 (Jansson and Farrar, 2012) models provide robust large-scale baselines, our adoption of Gaussian and shell-like parameterizations offers a mathematically tractable approach for representing localized plasma morphologies. This framework captures the dominant gradient variances essential for wave modulation, though it may smooth over fine-scale turbulent clumpiness observed in high-resolution maps (Leroy et al., 2013; Yamamoto et al., 2017).
@@
II.5 Quantitative Characterization of Galactic Profiles
To evaluate the location dependent KdV coefficients and the resulting KA soliton properties, we visualize how the background plasma parameters vary across our structured ISM model. Using the formulations in Sec. II.1II.4, we construct two-dimensional (2D) Galactic maps (Fig. 1) and detailed one-dimensional (1D) midplane profiles ($Z=0$; Figs. 2 and 3). These profiles capture the combined effects of the WIM and the embedded structures, serving as the environmental inputs for the soliton analysis presented in Sec. III. While the HII region’s impact is captured in the global 2D maps, we omit its corresponding 1D radial profiles for brevity. As these regions are modeled via simple Gaussian distributions for all plasma parameters, their 1D cross-sections lack the complex morphological features, such as shocked shells and interior plateaus, that characterize SWBs and SNRs. Consequently, our detailed 1D analysis focuses exclusively on the latter two structures to better illustrate the impact of sharp gradients and multi-component core-shell interactions on KA soliton dynamics.
For notational clarity in the derivation that follows, we adopt a shorthand in which the explicit dependence on galactic coordinates ($R,Z$) is dropped. Additionally, we replace the “total” subscript with a “0” subscript to denote equilibrium values at a specific location. Thus, we have the mappings: $n_{e,\text{total}}\rightarrow n_{e0}$, $n_{i,\text{total}}\rightarrow n_{i0}$, $B_{0,\text{total}}\rightarrow B_{0}$, and $T_{e,\text{total}}\rightarrow T_{e0}$. For the positron fraction, we simply adopt the notation $p_{\text{total}}\rightarrow p$. This convention streamlines the fluid equations while preserving their spatial dependence implicitly.
@@ II.5 Quantitative Characterization of Galactic Profiles
II.5.1 Galactic Background and H II Regions
@@
Fig. 1 presents the galactic 2D distribution of the plasma parameters in the $(R,Z)$ plane. All parameters are plotted on a $\log_{10}$ scale to capture the wide dynamic range of the ISM. The large scale background follows the disk-halo morphology, with magnetic field strength ($B_{0}$, Fig. 1a) and density ($n_{i0}$, Fig. 1b) decreasing gradually with Galactic height and radius. In the 2D maps, the background magnetic field appears around $\log_{10}(B_{0}/\mu\mathrm{G})\approx 0.85$, which corresponds to a physical field strength of $B_{0}\approx 7\,\mu\mathrm{G}$ , consistent with standard WIM values. Similarly, the background electron temperature (Fig. 1c) remains near $\log_{10}(T_{e0}/K)\approx 3.9$ ($T_{e0}\approx 8000$ K). The parameters adopted for the WIM background and the structural enhancements of the embedded astrophysical environments are given in Tables 1 and 2 respectively.
@@
$\sigma_{p,\text{scale}}$ 1.0 The Galactic positron source scale (kpc)
Embedded within this smooth background at $(R,Z)\approx(4,0)$ kpc is the H II region. In the 2D maps, it appears as a distinct, localized perturbation: a diamagnetic cavity in the magnetic field (Fig. 1a) coincident with a density enhancement (Fig. 1b) and a modest thermal bump (Fig. 1c). The high thermal pressure and reduced magnetic field within the H II region drive the local plasma beta significantly above unity ($\beta>1$). In Fig. 1d, this high-$\beta$ regime corresponds to the region masked in cyan. The implications of this $\beta$ distribution for soliton existence are analyzed in Sec. V.
@@
H II Region Electron density $n_{e,\mathrm{HII}}$ $\mathcal{G}$ $A_{n,\mathrm{HII}}=\texttt{2000 $\times$ $n_{e,\mathrm{bkg}}(R_{1},Z_{1})$}$, $\sigma_{n,\mathrm{HII}}=0.25$.

### v2

@@ 2.3.4 Model of Supernova Remnants
The expanding forward shock from an SNR sweeps up the surrounding interstellar gas, compressing the plasma into a dense, turbulent shell while simultaneously evacuating the interior (Chevalier, 1977; Vink, 2012). This process results in a complex density morphology consisting of a swept-up shell and a depleted cavity. In composite SNRs, this cavity also hosts a central PWN, injected by the relativistic wind of the compact remnant (Kennel and Coroniti, 1984; Gaensler and Slane, 2006). The shell is characterized by a strong peak in electron density. This enhancement is highly asymmetric, exhibiting a sharp density jump at the outer shock front followed by a gradual decline toward the interior due to post-shock relaxation (Reynolds, 1998; Orlando et al., 2007). The blast wave efficiently clears ambient ions from the remnant’s interior, creating a deep density cavity. Additionally, the PWN acts as an efficient source for electron-positron pairs, locally elevating the electron density in the core. Using the formalism from Sec. 2.2, we model this three-component morphology as a superposition of the PWN core, the cavity, and the shell:
@@
Since the PWN is a source of relativistic pairs, the positron density follows the same profile as electrons in the PWN (Kennel and Coroniti, 1984; Gaensler and Slane, 2006). We model this central positron source as a Gaussian distribution:
@@
In our SNR model, the magnetic field is structured in close analogy to the plasma density. The same forward shock that sweeps up and compresses the gas also compresses and amplifies the magnetic field lines via flux freezing (Vink, 2012). Consequently, the magnetic morphology exhibits a turbulent, high-field shell surrounding a comparatively weaker interior cavity that separates the shell from the central PWN core (Kennel and Coroniti, 1984; Gaensler and Slane, 2006). We model the total SNR magnetic field as the superposition of the shell and the PWN core:
@@
While the plasma density and magnetic field exhibit a hollow cavity structure, the temperature profile follows a distinctly different morphology. At the outer shock front, the kinetic energy of the sweeping ejecta is thermalized, creating an immediate jump in temperature (Chevalier, 1977; Vink, 2012). Interior to the shell, the physics of the Sedov-Taylor expansion phase dictates that while the density drops, the pressure remains high to support the expanding shell (Chevalier, 1977). Consequently, the temperature must rise towards the center ($T\propto P/n$) to maintain an approximately isobaric state. Furthermore, the center of the remnant is energized by the relativistic PWN (Gaensler and Slane, 2006; Kennel and Coroniti, 1984). To model this rising interior thermal structure, we treat the temperature as a superposition of shock heating at the shell and a broad, hot core filling the interior:
@@
2.4 Total ISM Profiles and Equilibrium Ion Density
@@
While the NE2001 (Cordes and Lazio, 2002) and JF12 (Jansson and Farrar, 2012) models provide robust large-scale baselines, our adoption of Gaussian and shell-like parameterizations offers a mathematically tractable approach for representing localized plasma morphologies. This framework captures the dominant gradient variances essential for wave modulation, though it may smooth over fine-scale turbulent clumpiness observed in high-resolution maps (Leroy et al., 2013; Yao et al., 2017).
@@
[…]
For notational clarity in the derivation and analysis that follows, we adopt a shorthand in which the explicit dependence on galactic coordinates ($R,Z$) is dropped. Additionally, we replace the “total” subscript with a “0” subscript to denote equilibrium values at a specific location. Thus, we have the mappings: $n_{e,\text{total}}\rightarrow n_{e0}$, $n_{i,\text{total}}\rightarrow n_{i0}$, $B_{0,\text{total}}\rightarrow B_{0}$, and $T_{e,\text{total}}\rightarrow T_{e0}$. For the positron fraction, we simply adopt the notation $p_{\text{total}}\rightarrow p$. This convention streamlines the fluid equations while preserving their spatial dependence implicitly.
2.5 Quantitative Characterization of Galactic Profiles
To evaluate the location dependent KdV coefficients and the resulting KA soliton properties, we visualize how the background plasma parameters vary across our structured ISM model. Using the formulations in Sec. 2.12.4, we construct two-dimensional (2D) Galactic maps (Fig. 1) and detailed one-dimensional (1D) midplane profiles ($Z=0$; Figs.2 and 3). These profiles capture the combined effects of the WIM and the embedded structures, serving as the environmental inputs for the soliton analysis presented in Sec. 3. While the HII region’s impact is captured in the global 2D maps, we omit its corresponding 1D radial profiles for brevity. As these regions are modeled via simple Gaussian distributions for all plasma parameters, their 1D cross-sections lack the complex morphological features, such as shocked shells and interior plateaus, that characterize SWBs and SNRs. Consequently, our detailed 1D analysis focuses exclusively on the latter two structures to better illustrate the impact of sharp gradients and multi-component core-shell interactions on KA soliton dynamics.
@@ 2.5 Quantitative Characterization of Galactic Profiles
2.5.1 Galactic Background and H II Regions
@@
Fig. 1 presents the galactic 2D distribution of the plasma parameters in the $(R,Z)$ plane. All parameters are plotted on a $\log_{10}$ scale to capture the wide dynamic range of the ISM. The large scale background follows the disk-halo morphology, with magnetic field strength ($B_{0}$, Fig. 1a) and density ($n_{i0}$, Fig. 1b) decreasing gradually with Galactic height and radius. In the 2D maps, the background magnetic field is maintained around $\log_{10}(B_{0}/\mu\mathrm{G})\approx 0.85$, which corresponds to a physical field strength of $B_{0}\approx 7\,\mu\mathrm{G}$ , consistent with standard WIM values. Similarly, the background electron temperature (Fig. 1c) remains near $\log_{10}(T_{e0}/K)\approx 3.9$ ($T_{e0}\approx 8000$ K). The parameters adopted for the WIM background and the structural enhancements of the embedded astrophysical environments are given in Tables 1 and 2 respectively.
Embedded within this smooth background at $(R,Z)\approx(4,0)$ kpc is the H II region. In the 2D maps, it appears as a distinct, localized perturbation: a diamagnetic cavity in the magnetic field (Fig. 1a) coincident with a density enhancement (Fig. 1b) and a modest thermal bump (Fig. 1c). The high thermal pressure and reduced magnetic field within the H II region drive the local plasma beta significantly above unity ($\beta>1$). In Fig. 1d, this high-$\beta$ regime corresponds to the region masked in cyan. The implications of this $\beta$ distribution for soliton existence are analyzed in Sec. 5.
2.5.2 Stellar-Wind Bubbles
[…]
2.5.3 Supernova Remnants
[…]
The SNR dominates the thermal map (Fig. 1c) with a massive temperature spike. The 1D profile (Fig. 3c) shows this temperature rising sharply at the shock front to $10^{6}$ K and remaining elevated throughout the interior, consistent with the Sedov-Taylor blast wave physics described in Sec. 2.3.4. Unlike the denser, collisional H II region where viscosity may damp the cascade, the shock heated plasma in the SNR shell is effectively collisionless, permitting the turbulent cascade to extend down to the kinetic scales required for KAW generation. Throughout this model, we assume the positrons are in thermal equilibrium with the electrons ($T_{p0}\approx T_{e0}$). Although the central PWNe is known to inject additional relativistic particle energy, this contribution is conservatively neglected in our model. PWNe thermal heating is typically confined to sub-parsec scales and is secondary to the broader, more intense shock heating at the shell, which dominates on the kpc-resolved grid (Temim et al., 2015). The distribution of plasma $\beta$ (Fig. 1d) at SNR exhibits a highly layered morphology. The shock heated shell, where $\beta>1$, appears as a cyan ring, while the deep interior core, where magnetic field is strong and $\beta<m_{e}/m_{i}$, is marked by a green dot. The detailed correspondence between these $\beta$ features and the allowable soliton regimes is discussed in Sec. 5.
3 Governing Fluid Equations for KA Solitons
Having established the large scale physical environment of the ISM, we now formulate the fluid equations that govern the dynamics of KA solitons. Our model assumes a collisionless, magnetized electron-positron-ion (e-p-i) plasma, consistent with conditions in the WIM. The dynamics are described within a local Cartesian frame $(x,z)$, where the equilibrium plasma parameters are determined by our multi-component ISM model at a specific galactic location $(R,Z)$. In this frame, positive ions are treated as a fluid, while the inertialess electrons and positrons follow superthermal kappa ($\kappa$) distributions, a common feature in space and astrophysical plasmas shaped by energetic processes.
To generalize the fluid equations, all physical quantities are normalized by characteristic scales that are themselves functions of the local ISM parameters. This ensures our results explicitly depend on the galactic location. Time is normalized by the ion-acoustic transit time across a gyroradius: $t=t^{\prime}\,V_{A}/\rho_{i}$; velocities by the Alfvén speed: $v_{ix,iz}=v_{ix,iz}^{\prime}/V_{A}$; potentials by characteristic energy scale $(k_{B}T_{e0}/e)$: $(\phi,\psi)=(\phi^{\prime},\psi^{\prime})/(k_{B}T_{e0}/e)$; spatial coordinates by the ion gyroradius: $x,z=(x^{\prime},z^{\prime})/\rho_{i}$; and densities by their local equilibrium values: $n_{j}=n_{j}^{\prime}/n_{j0}$. The key characteristic parameters are thus defined locally as: $V_{A}={B_{0}}/{\sqrt{4\pi n_{i0}m_{i}}}$, $\Omega_{i}={eB_{0}}/{m_{i}c}$, $C_{s}=\sqrt{{k_{B}T_{e0}}/{m_{i}}}$, $\rho_{i}={C_{s}}/{\Omega_{i}}$, $\beta={8\pi n_{i0}k_{B}T_{e0}}/{B_{0}^{2}}$. In this framework, the dynamics of KA solitons in a low-$\beta$ regime are governed by the dimensionless equations presented below (Eqs. 2024), where the dimensionless parameters explicitly depend on galactic location via our multi-component ISM model. This provides a direct path to studying KA solitons within specific, observable astronomical structures.
$v_{ix}=-\Lambda\frac{\partial^{2}\phi}{\partial x\partial t},$ (21)
where $\Lambda=\beta/2$. The equilibrium charge neutrality condition, expressed in terms of the density ratios as $\delta_{ei}=1+p,$ is determined explicitly by the local positron fraction $p$ (as defined in Eq. 17), where $\delta_{ei}\equiv n_{e0}/n_{i0}$. The normalized densities of the superthermal electrons and positrons, in response to the parallel potential $\psi$ of KA solitons, are given by (Singh et al., 2019):
$n_{l}=\left[1-{R_{l}}\,\frac{\psi}{\kappa_{l}-\tfrac{3}{2}}\right]^{-\kappa_{l}+\tfrac{1}{2}}\;\;\approx\;\;1+c_{1l}\psi+c_{2l}\psi^{2},$ (24)
where $l\in\{e,p\}$ denotes electrons ($e$) and positrons ($p$), with $R_{e}=+1,\quad R_{p}=-\alpha$, and the expansion coefficients are expressed as
$c_{1l}$ $=R_{l}\,\frac{\kappa_{l}-{1}/{2}}{\kappa_{l}-{3}/{2}},$ $c_{2l}$ $=R_{l}^{2}\,\frac{\kappa_{l}^{2}-{1}/{4}}{2(\kappa_{l}-{3}/{2})^{2}},$
where $\kappa_{l}$ denotes the spectral index quantifying the superthermality of species $l$, and $\alpha\equiv T_{e0}/T_{p0}$ represents the electron-to-positron temperature ratio. This set of self-consistent equations forms the basis for the derivation of the KdV equation in the subsequent sections.
4 Derivation of the KdV equation and its solution
The KdV equation is derived using the reductive perturbation method, and defining the independent stretched coordinate system $\xi$ and $\tau$ as:
where $\lambda$ is the phase velocity of the KAWs, $\epsilon$ is a small expansion parameter ($0<\epsilon\ll 1$) characterizing the weak nonlinearity. The geometric parameters $l_{x}$ and $l_{z}$ represent the direction cosines in the xzx-z plane, satisfying condition $l_{x}^{2}+l_{z}^{2}=1$. The dependent variables can be expanded as a power series in $\epsilon$:
where $S=(v_{ix},v_{iz},\psi)$. Note that while Eq. (19) defines the dimensional equilibrium ion density $n_{i0}$ used to construct the global maps, the variable $n_{i}$ in the expansions below refers to the dimensionless ion density normalized to this local background. Consequently, the equilibrium state corresponds to $n_{i}=1$ and $\psi=0$. According to plasma approximation, the normalized expression can be written as: $n_{i}=(1+p)\,n_{e}-p\,n_{p}.$ Using Eq. (24) in above equation, we obtain
$n_{i}=1+a_{1}\psi+a_{2}\psi^{2}+$... (27)
where $a_{1}=(1+p)\,c_{1e}+p\,c_{1p}$ and $a_{2}=(1+p)\,\delta_{ei}c_{2e}-p\,c_{2p}$. Using Eqs. (25), (26) and (27) in Eqs. (20)-(23) and comparing the coefficients of lowest powers of $\epsilon$, we obtain the first order governing Eqs. (36)-(39), given in Appendix C. On solving these first-order governing equations, we obtain the biquadratic dispersion equation
$a_{1}\lambda^{4}-(a_{1}+\beta)\,l_{z}^{2}\lambda^{2}+\beta l_{z}^{4}=0.$
On simplification of the above biquadratic equation, we obtain two different roots corresponding to KA mode and ion acoustic mode which are given respectively as: $\lambda^{2}=l_{z}^{2}$ and $\lambda^{2}={\beta l_{z}^{2}}/{a_{1}}$. The next order of $\epsilon$ yields the second order evolution Eqs. (40)-(43), given in Appendix D. Eliminating the second order perturbed quantities from these equations, the following KdV equation is obtained:
where $P=-a_{1}l_{z}$ is the nonlinear coefficient and $Q=-\frac{l_{x}^{2}l_{z}\beta}{2(a_{1}-\beta)}$ is dispersion coefficient. The stationary solution of the KdV equation (28) is obtained by introducing the traveling wave transformation $\eta=\xi-U\tau$. The resulting soliton solution is given by (Saini et al., 2015, 2017; Slathia et al., 2022):
$\psi^{(1)}=\psi_{0}\,\mathrm{sech}^{2}\!\left(\frac{\eta}{W}\right),$ (29)
5 Discussion
Figure 4: Two-dimensional Galactic maps of KA soliton (a) amplitude $\psi_{0}$ and (b) width W, at $\kappa_{e}=150,\kappa_{p}=100$, $\theta=85^{\circ}$, U=0.06U=-0.06 and $\sigma=1$. The parameters for the background (WIM) and the analytical profiles of embedded structures are given in Tables 1 and 2 respectively.
5.1 Amplitude and Width
The influence of electron superthermality on soliton morphology is shown in Figs. 5b,c and 6b,c. Solid green lines represent highly superthermal electrons ($\kappa_{e}=1.6$), dotted black lines represent moderate superthermality ($\kappa_{e}=2.5$), and dotted red lines represent near-Maxwellian distributions ($\kappa_{e}=150$). Across all environments, we observe a consistent trend where both the soliton amplitude and width increase as the plasma becomes more Maxwellian. This behavior is explained by the ability of the local electric potential to organize the plasma particles. In the superthermal regime where $\kappa_{e}$ is low, the plasma contains a large population of very high speed electrons in the distribution tail. Because these particles are so energetic, it is much harder for a local electric potential to trap them or hold them back to create a large charge separation. Since these particles can easily escape the potential well, the plasma cannot support a large potential hill. Therefore, the maximum amplitude ($\psi_{0}$) is naturally limited at a lower value compared to a Maxwellian plasma. In contrast, thermal particles are slower and easier to organize into a coherent structure. This allows the plasma to sustain a much higher potential drop before the structure becomes unstable, resulting in the higher amplitudes seen in the near-Maxwellian cases. This change in particle dynamics also regulates the soliton width. The excess energetic particles at low $\kappa_{e}$ enhance the nonlinear steepening of the wave. This strong nonlinearity acts to compress the soliton into a narrower and more compact structure. As the plasma thermalizes and $\kappa_{e}$ increases, this nonlinear compression effect weakens relative to the dispersive spreading. This allows the soliton to expand spatially, leading to the broader profiles observed in thermalized regions. This broadening becomes most prominent near the boundaries of the exclusion zones, where steep local gradients in temperature and density further increase the dispersive spreading, causing the soliton to spread before it reaches the stability threshold.
6 SUMMARY AND CONCLUSIONS
Acknowledgements.
Appendix A Model of H II Regions
Superposed on the quiescent background are distinct regions of significantly altered density, which we model as localized perturbations using specific analytical forms. First, H II regions represent strong positive electron density enhancements, $n_{e,\text{H\,II}}$. These are islands of plasma, often visible as bright emission nebulae, generated by intense photoionizing radiation from central O- or B-type massive stars (Haffner et al. 2009). Since these regions form from the gravitational collapse of dense molecular clouds, their internal density is significantly higher than the ambient ISM (Pagel et al. 1979; Dopita and Evans 1986; Brinchmann et al. 2008; Shirazi et al. 2014; Li et al. 2025). Using the analytical formalism for Localized structures introduced in the main text, Sec. 2.2, we model this contribution as a localized Gaussian enhancement:
In H II regions, the high thermal pressure of the hot, ionized gas drives an outward expansion. This expansion displaces the frozen-in field lines, creating a magnetic cavity with reduced field strength (Spitzer 1978; Ferrière 2001). We model this effect as a localized Gaussian depletion:
Inside H II regions, photoionization heating, where excess energy from ionizing photons is converted into kinetic energy of electrons, elevates the temperature above the background WIM (Draine 2011). Crucially, this temperature rise is limited by a thermostat effect, as temperature increases, cooling via collisionally excited forbidden lines rises exponentially, effectively limiting the equilibrium temperature typically near $10^{4}$ K (Osterbrock and Ferland 2006). We model this local thermal enhancement as:
Appendix B Model of Stellar-Wind Bubbles
The powerful radiatively driven winds from massive stars excavate large, low-density cavities in the ISM, surrounded by dense shells of swept-up gas (Castor et al. 1975; Weaver et al. 1977). To model this morphology, we construct the electron density profile as the superposition of a central depletion and a shell enhancement:
The magnetic structure of the bubble is governed by flux freezing. As the wind bubble inflates, the frozen-in magnetic field lines are stretched and diluted within the volume, while being strongly compressed into the dense shell at the boundary. We model this as:
Standard SWB models describe the interior as a region of hot, shocked stellar wind ($T\sim 10^{6}$ K) that maintains high pressure to support the bubble (Weaver et al. 1977). Because the sound speed is high, the interior evolves into an isobaric region with a nearly uniform temperature distribution. To reproduce this characteristic flat-top thermal profile, we model the interior temperature using the super-Gaussian plateau function defined in Eq. 4 in the main text. The total electron temperature at SWB includes this plateau and the shock-heated shell:
@@ Appendix B Model of Stellar-Wind Bubbles
$\sigma_{p,\text{scale}}$ 1.0 Galactic positron source scale (kpc)
@@
H II Region Electron density $n_{e,\mathrm{HII}}$ $\mathcal{G}$ $A_{n,\mathrm{HII}}=\text{2000 $\times$ $n_{e,\mathrm{bkg}}(R_{1},Z_{1})$}$, $\sigma_{n,\mathrm{HII}}=0.25$.